OpenStudy (anonymous):

The set {0, 1} is closed under which operation? (Points : 3) addition multiplication subtraction none of the above

OpenStudy (swissgirl):

A set is closed under a certain operation if by using that operation on all members in the set results in only members of the set Multiplication: 0*0=0 0*1=1 1*1=0 Addition: 0+0=0 0+1=1 1+1=2 Subtraction: 0-0=0 0-1=-1 1-1=0 1-0=1 So which operation would you think this set is closed for?

OpenStudy (anonymous):

multiplication

OpenStudy (swissgirl):

Yup

OpenStudy (anonymous):

could you help me with another?

OpenStudy (swissgirl):

Sure

OpenStudy (anonymous):

Which expression is greater than -|-0.4|? (Points : 3) |-(-0.4)| -(-(-0.4)) -|-(-0.4)| -|-0.4|

OpenStudy (swissgirl):

Ok lets start with -|-.4| -|-.4|=-.4

OpenStudy (anonymous):

i know that -|-.4| is -.4

OpenStudy (swissgirl):

|-(-.4)|=|.4|=.4

OpenStudy (anonymous):

the next one would be -.4 right @swissgirl